OpenStudy (anonymous):

Probability and Statistics question. Anyone knows a proof of this or at least a book with a proof of this? (Second post in LaTeX)

OpenStudy (anonymous):

Difference between two means with unknow variances. Estimator: \[\frac{\bar{x}-\bar{y}-\Delta_0}{Sp \sqrt{\frac{1}{n}+\frac{1}{m}}}\] With \[Sp^2 = \frac{s^2_x(n-1)+s^2_y(m-1)}{n+m-2}\] I don't know why but I can't put "S" in latex without it changing to: \[\S\] So I used "s" to x and y samples variations.

OpenStudy (anonymous):

Nope. Sorry :(